Isopropyl alcohol is mixed with water to produce a 39.0 % (v/v) alcohol solution. How many milliliters of each component are present in 855 mL of this solution? Assume that volumes are additive. alcohol: mL water: mL
Answer #1

% v/v = (volume of alcohol / total volume) x 100

% v/v = 39.0

volume of alcohol = ?

total volume = 855 mL

39.0 = (volume of alcohol / 855) x 100

volume of alcohol / 855 = 0.39

volume of alcohol = 333.45 mL

volume of water = 855 - 333.45

volume of water = 521.55 mL